# V2XPnP: Vehicle-to-Everything Spatio-Temporal Fusion for Multi-Agent Perception and Prediction

### v1

@@ Abstract
Vehicle-to-everything (V2X) technologies offer a promising paradigm to mitigate the limitations of constrained observability in single-vehicle systems. Prior work primarily focuses on single-frame cooperative perception, which fuses agents’ information across different spatial locations but ignores temporal cues and temporal tasks (e.g., temporal perception and prediction). In this paper, we focus on temporal perception and prediction tasks in V2X scenarios and design one-step and multi-step communication strategies (when to transmit) as well as examine their integration with three fusion strategies - early, late, and intermediate (what to transmit), providing comprehensive benchmarks with various fusion models (how to fuse). Furthermore, we propose V2XPnP, a novel intermediate fusion framework within one-step communication for end-to-end perception and prediction. Our framework employs a unified Transformer-based architecture to effectively model complex spatiotemporal relationships across temporal per-frame, spatial per-agent, and high-definition map. Moreover, we introduce the V2XPnP Sequential Dataset that supports all V2X cooperation modes and addresses the limitations of existing real-world datasets, which are restricted to single-frame or single-mode cooperation. Extensive experiments demonstrate our framework outperforms state-of-the-art methods in both perception and prediction tasks.
footnotetext: Corresponding author {zhiyuh,jiaqima}@ucla.edu
@@ 1 Introduction
Despite their potential, existing works focus on frame-by-frame cooperative detection [41, 37, 51], which aggregates information from agents at different spatial locations. However, these works overlook temporal cues across sequential frames, which are important for locating previously visible but currently undetected objects [44] and predicting objects’ future trajectories [29]. The potential of V2X for enhancing these temporal tasks, particularly in cooperative temporal perception and prediction, remains largely unexplored. We aim to address critical questions in multi-agent cooperation: (1) What information to transmit? (2) When to transmit it? (3) How to fuse information across spatial and temporal dimensions? To address what to transmit, we expand the three fusion strategies in single-frame cooperative perception (i.e., early, late, and intermediate) to incorporate the temporal dimension. Regarding when to transmit within the temporal dimension, we introduce one-step and multi-step communication strategies to capture multi-frame temporal information. For how to fuse, we conduct a systematic analysis across various spatio-temporal fusion strategies, providing comprehensive benchmarks for cooperative perception and prediction tasks across all V2X collaboration modes.
Among these strategies, we advocate intermediate fusion within one-step communication, because it effectively balances the trade-off between accuracy and increased transmission load. Moreover, its capability to transmit intermediate spatial-temporal features makes it well-suited for end-to-end perception and prediction, allowing for feature sharing across multiple tasks and reducing computational demands, as shown in Fig. 1. Based on this strategy, we propose V2XPnP, a V2X spatio-temporal fusion framework leveraging a unified Transformer structure for effective spatial-temporal fusion, encompassing temporal attention, self-spatial attention, multi-agent spatial attention, and map attention. Each agent first extracts its inter-frame and self-spatial features, which can support single-vehicle perception and prediction while reducing the communication load, and then the multi-agent spatial attention model fuses the single-agent feature across different agents.
@@
    We conduct extensive analysis across various spatio-temporal fusion strategies and provide comprehensive benchmarks for cooperative perception and prediction in all V2X collaboration modes, demonstrating the state-of-the-art performance of the proposed model in cooperative temporal perception and prediction tasks.
@@ 2 Related Work
End-to-end Perception and Prediction. Safe autonomous driving fundamentally depends on accurate perception and prediction [17, 25, 39, 47]. In single-agent systems, significant efforts have been dedicated to temporal perception and prediction [26, 17], with numerous end-to-end perception and prediction frameworks emerging [29, 27, 11]. These frameworks offer computational efficiency by sharing information across tasks and mitigating error propagation inherent in modular architectures. For instance, FaF [29] employed two strategies to extract the temporal features and directly output the results. PnPNet [27] incorporated a loop tracking module to improve performance. Recently, Bird’s-Eye-View (BEV)-based approaches have gained prominence, focusing on leveraging camera data [11, 14].
V2X Perception and Prediction. Extensive studies have explored V2X perception, with intermediate fusion as a widely adopted approach. OPV2V [42] proposed an attention-based fusion pipeline, while V2X-ViT [41] and CoBEVT [40] utilized vision Transformer to fuse spatial features. However, coordinating a multi-agent system for temporal tasks remains an open challenge. In the prediction domain, extensive exploration of deep learning models, from CNN and LSTM [8] to Transformer [33] and graph network [19], has been conducted to model inter-agent interactions. However, the limited short-term visibility of individual vehicles continues to restrict prediction accuracy. Cooperative prediction leveraging V2X has shown potential, though research remains preliminary [46, 36, 44]. For instance, trajectory-level fusion has demonstrated improved prediction accuracy compared to no fusion [30]. Moreover, V2Vnet [35] utilized a graph neural network to perform joint perception and prediction. Nevertheless, the field currently lacks a comprehensive framework for V2X spatio-temporal tasks, and further research is still needed to fully realize the advantages of perception and prediction offered by V2X systems.
Real-world Driving Datasets. Public driving datasets have been instrumental in advancing autonomous driving research. Early sequential datasets, such as NGSIM [1] and HighD [21], provide object trajectories on highways [50], but lack perception data. The following datasets - nuScenes [4], Argoverse [5], and Waymo [34] - collected real perception data with AVs and extended to urban environments. However, these datasets can not support V2X research due to the single-agent view. To address this, simulated V2X datasets, such as V2XSet [41], were developed. Recently, datasets including Dair-V2X [45], V2V4Real [43], RCooper [13], and V2X-Real [38] contributed significant advancements in real-world data across V2I, V2V, I2I, and V2X perception tasks. However, the lack of real-world sequential V2X datasets remains. The only work in this field, V2X-Seq [46], only provides sequential V2I data with download restrictions.
@@ 3.2 V2XPnP Framework
Map Feature Extraction. The HD map is directly accessed for each agent without V2X fusion. We project the vectorized HD map to BEV space by incorporating the map polylines into each BEV feature grid. We first employ a multi-layer perceptron ($\operatorname{MLP}$) to encode the surrounding map polylines for each grid $\mathbf{M}\in\mathbb{R}^{H\times W\times N_{m}\times n\times D}$, resulting in the map feature $\mathbf{F}_{m}\in\mathbb{R}^{H\times W\times N_{m}\times C}$. Here, $N_{m}$ and $n$ represent the number of map polylines and the waypoints per polyline, while $D$ represents waypoint attributes (i.e., position and lane type). The map encoding is expressed as:
$\mathbf{F}_{m}=\phi\left(\operatorname{MLP}\left(\mathbf{M}\right)\right),$ (1)
where $\phi$ denotes max-pooling on the waypoint axis. Then, a map-BEV attention is introduced to inject the map feature into the BEV feature. We concatenate the BEV and map feature $\mathbf{F}_{bm}=[\mathbf{F}^{{}^{\prime}},\mathbf{F}_{m}]\in\mathbb{R}^{H\times W\times(1+N_{m})\times C}$, and add the position embedding $\mathbf{P}_{m}$ based on sinusoidal positional encoding. The final content feature map $\mathbf{F}\in\mathbb{R}^{H\times W\times C}$ is updated by multi-head self-attention ($\operatorname{MHSA}$) as follows:
@@ 3.3 Spatio-Temporal Fusion Transformer
Multi-agent Spatial Attention. This block facilitates inter-agent fusion by aggregating BEV feature maps across various spatial positions from multiple agents. Considering the different deployment positions and capabilities of vehicle and infrastructure sensors, the multi-agent spatial Transformer is heterogeneous with individual learnable weights for different interaction pairs (i.e., V-I, V-V, I-V, and I-I). The attention token of $i$ th agent $\mathbf{F}_{i,m}^{sp}$ is modulated by its type $m$ embedding, and is weighted by the relation matrix $\mathbf{W}_{\text{att}}^{(e_{i,j})}$ between edge $e_{i,j}$ during aggregation with agent $j$ of type $n$:
@@ 3.4 Learning Objective
The learning objective is comprised of temporal perception and prediction tasks. First, we define the perception loss as the combination of regression $\mathcal{L}_{\text{reg}}$ and classification loss $\mathcal{L}_{\text{cla}}$ of the predefined anchor box. Specifically, the smooth $\ell_{1}$-loss is leveraged for the regression part, and the focal loss [28] is utilized for classification. Second, we define the prediction loss $\mathcal{L}_{\text{pred }}$ as $\ell_{2}$-loss between the prediction points sequence with ground truth trajectory. The final loss function is their weighted sum:
$\mathcal{L}=w_{\text{reg }}\mathcal{L}_{\text{reg }}+w_{\text{cla}}\mathcal{L}_{\text{cla }}+w_{\text{pred }}\mathcal{L}_{\text{pred }},$ (5)
where $w_{\text{reg }}$, $w_{\text{cla}}$, and $w_{\text{pred }}$ represent balance weights.
@@ 4 V2XPnP Sequential Dataset
We introduce the V2XPnP-Sequential dataset, the first large-scale, real-world V2X sequential dataset featuring multiple agents and all collaboration modes. This dataset comprises 100 scenarios, each spanning 95 to 283 frames captured at 10 Hz. The dataset comprises two data sequences from CAV perception (point clouds and camera images) and two data sequences from infrastructure perception, as shown in Fig. 4(b). We also provide corresponding vector maps and point-cloud maps for all collection areas, as shown in Fig. 4(c). Ten object categories are included, and the average trajectory length and frequency of each category are shown in Fig. 4(d). Further details on data annotation, trajectory and map generation are provided in the supplementary.
@@ 4.1 Data Acquisition
V2X temporal tasks require diverse time-consistent perception data and object behavior data. We choose urban arterial roads, expressways, and intersections to collect V2X data. The sensor configurations for the two CAVs and two infrastructures are shown in Fig. 4(a). By permuting and combining the behavior patterns of CAVs (such as overtaking, platooning, turning, etc.), we designed a total of about 60 interaction pairs between CAVs. From more than 66h driving logs, we annotate 32 representative scenarios and process data from the non-sequential V2X-Real [38] with a V2X sequential data processing pipeline, detailed in Sec. 4.2. Since perception tasks do not need to consider the continuity of object labeling, identical objects in V2X-Real may be assigned different IDs within a sequence, leading to data fragmentation and limiting support for temporal tasks. The final dataset is yielded by processing our collected data and V2X-Real data.
@@ 5.1 Experimental Setup
Evaluation Metrics. Following the detection evaluation protocol in [38], we measure detection performance using Average Precision (AP) at an IoU threshold of 0.5. For prediction, we provide the results of commonly used metrics [32, 10, 31, 3], including Average Displacement Error (ADE), Final Displacement Error (FDE), and Miss Rate (MR) within a 2-meter threshold. However, prediction accuracy is influenced by false positives and missed detections from the perception module. For example, a poorly performing perception module that detects only a simple object in a straight-line trajectory can misleadingly inflate the prediction accuracy. To address this, we employ the End-to-end Perception and Prediction Accuracy (EPA) metric [11] to jointly evaluate perception and prediction performance.
$EPA=\frac{|\hat{S}_{\text{match, hit}}|-\alpha N_{\text{FP}}}{N_{\text{GT}}},$ (6)
@@
V2VNet* [35] 48.6 2.10 3.75 42.3 25.3
F-Cooper* [6] 66.0 1.35 2.56 36.1 38.7
V2XPnP (Ours) 71.6 +5.6 1.35 2.36 31.7 48.2 +9.5
@@
V2VNet* [35] 33.6 1.95 3.53 44.2 16.3
F-Cooper* [6] 60.2 1.21 2.32 36.3 36.3
V2XPnP (Ours) 71.0 +10.8 1.18 2.16 34.0 46.0 +9.7
@@
V2VNet* [35] 43.1 3.10 5.55 46.8 19.4
F-Cooper* [6] 60.2 1.69 3.22 41.1 34.4
V2XPnP (Ours) 70.5 +10.3 1.78 3.28 39.9 40.6 +6.2
@@
V2VNet* [35] 41.1 1.83 3.34 40.4 23.2
F-Cooper* [6] 58.6 1.34 2.58 40.0 33.6
V2XPnP (Ours) 69.2 +10.6 1.26 2.31 36.5 42.8 +9.2
Collaboration Modes. The V2XPnP Sequential Dataset supports various V2X collaboration modes by organizing data based on specific interaction patterns. Vehicle-Centric (VC): The ego CAV is the focal agent, communicating with other CAVs and infrastructure. Infrastructure-Centric (IC): Infrastructure is the central entity, communicating with other infrastructures and CAVs. Vehicle-to-Vehicle (V2V): The ego CAV communicates exclusively with other CAVs without involving infrastructure. Infrastructure-to-Infrastructure (I2I): Infrastructure units share data only with each other.
Implementation Details. During the testing stage, we select a fixed agent as the ego agent in each cooperative scenario, while the ego agent is shuffled and randomly selected during training. Following the real-time testing and evaluation [48], we set the communication range for the ego agent to 50 meters and evaluate surrounding agents within a range of $x\in[-70,70]$m and $y\in[-40,40]$m. Messages beyond 50 meters are discarded by the ego agent. Besides, the history length is 2s (2Hz), and the prediction horizon is 3s (2Hz). The train/validation/test data splits are 80/6/14 scenarios. All models are trained using the Adam optimizer [20] with early stopping on NVIDIA L40S GPUs. Additional training and model details are provided in the supplementary materials.

### v2

@@ Abstract
Vehicle-to-everything (V2X) technologies offer a promising paradigm to mitigate the limitations of constrained observability in single-vehicle systems. Prior work primarily focuses on single-frame cooperative perception, which fuses agents’ information across different spatial locations but ignores temporal cues and temporal tasks (e.g., temporal perception and prediction). In this paper, we focus on the spatio-temporal fusion in V2X scenarios and design one-step and multi-step communication strategies (when to transmit) as well as examine their integration with three fusion strategies - early, late, and intermediate (what to transmit), providing comprehensive benchmarks with 11 fusion models (how to fuse). Furthermore, we propose V2XPnP, a novel intermediate fusion framework within one-step communication for end-to-end perception and prediction. Our framework employs a unified Transformer-based architecture to effectively model complex spatio-temporal relationships across multiple agents, frames, and high-definition map. Moreover, we introduce the V2XPnP Sequential Dataset that supports all V2X collaboration modes and addresses the limitations of existing real-world datasets, which are restricted to single-frame or single-mode cooperation. Extensive experiments demonstrate our framework outperforms state-of-the-art methods in both perception and prediction tasks. The codebase and dataset will be released to facilitate future V2X research.
@@ 1 Introduction
Despite their potential, existing works focus on frame-by-frame cooperative detection [32, 43, 57, 26], which aggregates information from agents at different spatial locations. However, these works overlook temporal cues across sequential frames, which are important for locating previously visible but currently undetected objects [48] and predicting object future trajectories [33]. Although some work [51, 42] incorporate short-term temporal cues (0.5 s) in single-frame perception to mitigate asynchrony, the broader challenge of efficiently aggregating multi-agent and multi-frame information and supporting long-term temporal tasks, such as motion prediction, remains largely unexplored. Therefore, we aim to address critical questions in multi-agent multi-frame cooperation: (1) What information to transmit? (2) When to transmit it? (3) How to fuse information across multi-agent spatial and temporal dimensions? To address what to transmit, we expand the three fusion strategies in single-frame cooperative perception (i.e., early, late, and intermediate) to incorporate the temporal dimension. Regarding when to transmit, we introduce one-step and multi-step communication strategies to capture multi-frame temporal information. For how to fuse, we conduct a systematic analysis across various spatio-temporal fusion strategies, providing comprehensive benchmarks for cooperative perception and prediction tasks across all V2X collaboration modes.
Among these strategies, we advocate intermediate fusion within one-step communication, because it effectively balances the trade-off between accuracy and increased transmission load. Moreover, its capability to transmit intermediate spatial-temporal features makes it well-suited for end-to-end perception and prediction, allowing for feature sharing across multiple tasks, as shown in Fig. 1. Based on this strategy, we propose V2XPnP, a V2X spatio-temporal fusion framework leveraging a unified Transformer structure for effective spatial-temporal fusion, encompassing temporal attention, self-spatial attention, multi-agent spatial attention, and map attention. Each agent first extracts its inter-frame and self-spatial features, which can support single-vehicle perception and prediction while reducing the communication load, and then the multi-agent spatial attention model fuses the single-agent feature across different agents.
@@
    We conduct extensive analysis across various spatio-temporal fusion strategies and benchmarks 11 fusion models for cooperative perception and prediction in all V2X collaboration modes, demonstrating the state-of-the-art performance of the proposed model.
@@ 2 Related Work
End-to-end Perception and Prediction. Safe autonomous driving fundamentally depends on accurate perception and prediction [18, 54]. In single-agent systems, significant efforts have been dedicated to temporal perception and prediction [29], leading to the development of various end-to-end frameworks. These frameworks enhance computational efficiency by sharing information across tasks and mitigating error propagation inherent in modular architectures. FaF [33] and PnPNet [30] focus on Lidar-based joint perception and prediction, while occupancy flow methods [3, 2] provide detailed spatial-temporal information. Recently, Bird’s-Eye-View (BEV)-based approaches with camera data have gained prominence [12, 16, 21].
V2X Perception and Prediction. V2X perception has been extensively explored, with intermediate fusion emerging as a widely adopted strategy [46, 45, 10]. FFNet [51] and CoBEVFlow [42] utilize historical information (0.5s) from collaborators to mitigate asynchrony, and SCOPE integrates ego-history for detection. However, coordinating multi-agent systems for long-term temporal tasks remains an open challenge. In the prediction domain, deep learning models have been extensively studied for modeling inter-agent interactions [37, 20]. However, the limited short-term visibility of individual vehicles continues to restrict prediction accuracy. Cooperative prediction leveraging V2X has shown potential, though research remains preliminary [53, 41, 48, 34]. To integrate perception and prediction within an end-to-end framework, V2VNet [40] employs a graph neural network for spatio-temporal fusion. UniV2X [52] extended the end-to-end system to support downstream tasks, but with a simplified spatio-temporal fusion module. Despite these advancements, a comprehensive framework for V2X-based spatio-temporal fusion is still lacking.
Real-world Driving Datasets. Public datasets have been instrumental in advancing autonomous driving research. Early sequential datasets [1, 23] provided only object trajectories on highways [56], but lack perception data. The following datasets, such as nuScenes [5] and Waymo [39], introduced real-world urban data but were limited to single-agent perspectives, rendering them unsuitable for V2X research. Thus, simulated datasets like V2XSet [45] were developed. Recently, datasets including Dair-V2X [49], V2V4Real [47], RCooper [14], and V2X-Real [44] significant contributed to real-world data in the V2I, V2V, I2I and V2X modes. However, real-world sequential V2X datasets covering all collaboration modes remain scarce. V2X-Seq [50] is the only sequential dataset incorporated with various behavior and map data for prediction tasks; however, is limited to V2I data and has restricted accessibility with download constraints.
@@ 3.2 V2XPnP Framework
Map Feature Extraction. The HD map is directly accessed for each agent without V2X fusion. We project the vectorized HD map to BEV space by incorporating the map polylines into each BEV feature grid. We first employ a multi-layer perceptron ($\operatorname{MLP}$) to encode the surrounding map polylines for each grid $\mathbf{M}\in\mathbb{R}^{H\times W\times N_{m}\times n\times D}$, resulting in the map feature $\mathbf{F}_{m}\in\mathbb{R}^{H\times W\times N_{m}\times C}$. Here, $N_{m}$ and $n$ represent the number of map polylines and the waypoints per polyline, while $D$ represents waypoint attributes (i.e., position and lane type). The map encoding is expressed as: $\mathbf{F}_{m}=\phi\left(\operatorname{MLP}\left(\mathbf{M}\right)\right)$, where $\phi$ denotes max-pooling on the waypoint axis. Then, a map-BEV attention is introduced to inject the map feature into the BEV feature. We concatenate the BEV and map feature $\mathbf{F}_{bm}=[\mathbf{F}^{{}^{\prime}},\mathbf{F}_{m}]\in\mathbb{R}^{H\times W\times(1+N_{m})\times C}$, and add the position embedding $\mathbf{P}_{m}$ based on sinusoidal positional encoding. The final content feature map $\mathbf{F}\in\mathbb{R}^{H\times W\times C}$ is updated by multi-head self-attention ($\operatorname{MHSA}$) as follows:
@@ 3.3 Spatio-Temporal Fusion Transformer
Multi-agent Spatial Attention. This block facilitates inter-agent fusion by aggregating BEV feature maps from multiple agents. Considering the different deployment positions and capabilities of vehicle and infrastructure sensors, the multi-agent spatial Transformer is heterogeneous with individual learnable weights for different interaction pairs (i.e., V-I, V-V, I-V, and I-I). The attention token of $i$ th agent $\mathbf{F}_{i,m}^{sp}$ is modulated by its type $m$ embedding, and is weighted by the relation matrix $\mathbf{W}_{\text{att}}^{(e_{i,j})}$ between edge $e_{i,j}$ during aggregation with agent $j$ of type $n$:
@@ 3.4 Learning Objective
The learning objective is comprised of temporal perception and prediction tasks. First, we define the perception loss as the combination of regression $\mathcal{L}_{\text{reg}}$ and classification loss $\mathcal{L}_{\text{cla}}$ of the predefined anchor box. Specifically, the smooth $\ell_{1}$-loss is leveraged for the regression part, and the focal loss [31] is utilized for classification. Second, we define the prediction loss $\mathcal{L}_{\text{pred }}$ as $\ell_{2}$-loss between the prediction points sequence with ground truth trajectory. The final loss function is the weighted sum of $\mathcal{L}_{\text{reg }},\mathcal{L}_{\text{cla }}$, and $\mathcal{L}_{\text{pred }}$.
@@ 4 V2XPnP Sequential Dataset
We introduce the V2XPnP-Sequential dataset, the first large-scale, real-world V2X sequential dataset featuring multiple agents and all collaboration modes. This dataset comprises 100 scenarios (49 2V+2I scenarios, 42 V2V scenarios, and 9 V+2I scenarios), each spanning 95 to 283 frames captured at 10 Hz. The dataset comprises two data sequences from CAV perception (point clouds and camera images) and two data sequences from infrastructure perception, as shown in Fig. 4(b). We also provide corresponding vector maps and point-cloud maps for all collection areas, as shown in Fig. 4(c). Ten object categories are included, and the average trajectory length and frequency of each category are shown in Fig. 4(d). Further details on data visualization, annotation, trajectory and map generation are provided in the supplementary.
@@ 4.1 Data Acquisition
V2X temporal tasks require diverse time-consistent perception data and object behavior data. We choose urban arterial roads, expressways, and intersections to collect V2X data. The sensor configurations for two CAVs and two infrastructures are shown in Fig. 4(a). By permuting and combining the behavior patterns of CAVs (such as overtaking, platooning, turning, etc.), we designed a total of about 60 interaction pairs between CAVs to collect data. From more than 66h driving logs, we annotate 32 representative scenarios. We also process the non-sequential data from V2X-Real [44] with a V2X sequential data processing pipeline, detailed in Sec. 4.2. Since perception tasks do not need to consider labeling consistency, identical objects may be assigned different IDs within a sequence, leading to data fragmentation and limiting support for temporal tasks. The final dataset is yielded by processing our collected data and V2X-Real data.
@@ 5.1 Experimental Setup
Evaluation Metrics. Following the detection evaluation protocol in [44], we measure detection performance using Average Precision (AP) at an IoU threshold of 0.5. For prediction, we provide the results of commonly used metrics [36, 11, 35], including Average Displacement Error (ADE), Final Displacement Error (FDE), and Miss Rate (MR) within a 2-meter threshold. However, prediction accuracy is influenced by false positives and missed detections from the perception module. For example, a poorly performing perception module that detects only a simple object in a straight-line trajectory can misleadingly inflate the prediction accuracy. To address this, we employ the End-to-end Perception and Prediction Accuracy (EPA) metric [12] to jointly evaluate perception and prediction performance.
$\text{EPA}=\frac{|\hat{S}_{\text{match, hit}}|-\alpha N_{\text{FP}}}{N_{\text{GT}}},$ (4)
@@
CoBEVFlow* [42] 63.3 1.36 2.49 33.0 41.9
FFNet* [51] 64.6 1.36 2.47 34.7 42.3
V2X-ViT* [45] 69.6 1.39 2.56 35.2 44.7
V2XPnP (Ours) 71.6 +2.0 1.35 2.36 31.7 48.2 +3.5
@@
CoBEVFlow* [42] 57.6 1.38 2.58 31.0 32.5
FFNet* [51] 61.0 1.18 2.18 35.1 37.5
V2X-ViT* [45] 69.3 1.27 2.39 35.4 43.3
V2XPnP (Ours) 71.0 +1.7 1.18 2.16 34.0 46.0 +2.7
@@
CoBEVFlow* [42] 58.7 1.72 3.15 40.3 33.6
FFNet* [51] 56.5 1.68 3.12 39.8 31.2
V2X-ViT* [45] 64.6 1.68 3.13 39.8 36.7
V2XPnP (Ours) 70.5 +5.9 1.78 3.28 39.9 40.6 +3.9
@@
CoBEVFlow* [42] 58.4 1.31 2.61 41.5 33.0
FFNet* [51] 66.1 1.41 2.59 36.3 40.9
V2X-ViT* [45] 65.4 1.22 2.33 35.9 41.3
V2XPnP (Ours) 69.2 +3.1 1.26 2.31 36.5 42.8 +1.2
Collaboration Modes. The V2XPnP Sequential Dataset supports various V2X collaboration modes by organizing data with specific interaction patterns. Vehicle-Centric (VC): The ego CAV is the focal agent, communicating with other CAVs and infrastructure (Infra). Infrastructure-Centric (IC): Infrastructure is the central entity, communicating with other Infras and CAVs. Vehicle-to-Vehicle (V2V): The ego CAV communicates exclusively with other CAVs without involving Infra. Infrastructure-to-Infrastructure (I2I): Infra shares data only with each other. Each VC and IC scenario includes 2-4 agents, which is close to real-world V2X settings and can evaluate model generalization across diverse V2X scenarios, whereas each V2V and I2I scenario has two agents.
Implementation Details. During the testing stage, we select a fixed agent as the ego agent in each cooperative scenario, while the ego agent is shuffled and randomly selected during training. Following the real-time setting [55], we set the communication range to 50 meters and evaluate surrounding agents within a range of $x\in[-70,70]$m and $y\in[-40,40]$m. Messages beyond 50 meters are discarded. Besides, the history length is 2s (2Hz), and the prediction horizon is 3s (2Hz). The train/validation/test data splits are 80/6/14 scenarios. Additional training and model details are provided in the supplementary materials.